# MGDA Converges under Generalized Smoothness, Provably

### v4

@@ 2.2 Pareto concepts in multi-objective optimization (MOO)
As described before, MOO aims to find points at which there is no common descent direction for all objectives. Considering two points $x_{1},x_{2}\in\mathbb{R}^{m}$, we claim that $x_{1}$ dominates $x_{2}$ if $f_{i}(x_{1})\geq f_{i}(x_{2})$ for all $i\in[K]$ and $F(x_{1})\neq F(x_{2})$. We say a point is Pareto optimal if it is not dominated by any other point. In other words, we cannot improve one objective without compromising another when we reach a Pareto optimal point. In the general non-convex setting, MOO aims to find a Pareto stationary point defined as follows.
@@
We say $x\in\mathbb{R}^{m}$ is a Pareto stationary point if $\min_{w\in\mathcal{W}}\|\nabla F(x)w\|^{2}=0$. In practice, we call $x$ an $\epsilon$-accurate Pareto stationary point if $\min_{w\in\mathcal{W}}\|\nabla F(x)w\|^{2}\leq\epsilon^{2}$.
@@ 2.3 Existing MOO Algorithms
Deterministic MOO. One of the big challenges of MOO is the gradient conflict, i.e., the gradients of different objectives may vary heavily in scale such that the largest gradient dominates the update direction. As a result, the performance of those objectives with smaller gradients [37] may be significantly compromised. As the most fundamental MOO algorithm, MGDA tends to find a balanced update direction for all objectives by considering the minimum improvement across all objectives and maximize it by solving the following problem
@@
where $\mathcal{W}$ is the probability simplex over $[K]$, and the regularization term $-\frac{1}{2}\|d\|^{2}$ is to regulate the magnitude of our update direction. The solution to the problem in eq. 3 can be obtained by solving the following problem [12]:
@@
The above approach has been widely used in various variants of MGDA such as SDMGrad, CAGrad, and PCGrad [35, 37, 25].
Stochastic MOO. SMG [27] is the first stochastic MGDA. It directly replaces the gradients with stochastic gradients and the update rule becomes
@@
where $\nabla F(x;s)$ is the estimate of $\nabla F(x)$ based on the sample $s$. However, this leads to a biased gradient estimation of the update direction $d_{s}^{*}$, and thus it requires an increasing batch size. To solve this issue, another work MoCo [13] introduces a tracking variable $Y$ as a stochastic estimation of the true gradient. Afterward, a double-sampling strategy is proposed by [6, 35] to generate a near-unbiased update direction.
Strong assumptions in analysis. All the works mentioned above require bounded gradients such as [6, 13, 35] or $L$-smoothness such as [3, 25, 29, 36]. Their analyses do not apply to generalized $\ell$-smoothness objectives, since the Lipschitz constant is potentially infinity.
@@ 2.4 Conflict-avoidant (CA) direction and CA distance
The larger the CA distance is, the further the estimated update direction will be away from the CA direction, and the more conflict there will be. In single-loop algorithms, MoCo and MoDo [13, 6], the average CA distance over iteration is of the order of $\epsilon$ while the double-loop algorithm SDMGrad [35] guarantees an $\epsilon$-order CA distance in every iteration. In this work, we analyze the CA distance in both cases and provide convergence results.
@@ 3.1 MGDA with and without Warm Start
Besides the benefit of a unique solution, adding an $\ell_{2}$ regularization term also makes $w_{\rho}^{*}(x)$ Lipschitz continuous [13]. Note that $w^{*}(x)$ may not be Lipschitz continuous because $\nabla F(x)^{\top}\nabla F(x)$ may not be positive definite. Nevertheless, the analysis of CA distance is difficult because $w^{*}$ may not be Lipschitz continuous. Thus, we will characterize the gap between $w^{*}$ and $w_{\rho}^{*}$ plus the change of $w_{\rho}^{*}$ after adding this $\ell_{2}$ regularization term. As a result, the update rules become Lines 5-6 in Algorithm 1. We first update $w_{t}$ by a projected gradient descent process and compute the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to update model parameters.
@@ 3.2 Stochastic MGDA with Double Sampling
In the stochastic setting, our algorithm keeps the same structure, having a warm start process if we aim to control the CA distance in every iteration. In Algorithm 2, we do the same projected gradient descent without using stochastic gradients. This is because we only need to compute $\nabla F(x_{0})^{\top}\nabla F(x_{0})$ once and reuse it in the whole loop, which does not bring a computational burden. Then in the update loop, we update the weight and model parameters accordingly. We use a double-sampling strategy here to make the weight gradient estimator unbiased [35] such that $d_{t}$ is a near-unbiased multi-gradient $\mathbb{E}[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]=\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t},$ where $\nabla G_{2}(x_{t})$ and $\nabla G_{3}(x_{t})$ are independent and unbiased estimates of $\nabla F(x_{t})$. Similarly, we do not involve a warm start process if we require the average CA distance to be small.
@@
3.3 MGDA with Fast Approximation
It can be seen from Algorithm 1 and Algorithm 3 that MGDA requires $\mathcal{O}(K)$ space and time to compute and store all task gradients at each iteration for updating the weight $w_{t}$. This becomes a drawback when the number of tasks or the model size is large. Motivated by [24], one solution is to use the Taylor Theorem to approximate the gradient for updating the weight $w_{t}$ as
where $R(x_{t})$ is the remainder term and it takes the order $R(x_{t})=o(\|x_{t}-x_{t+1}\|^{2})$, which can be made sufficiently small by adjusting the step size. By incorporating this fast approximation (FA) in Algorithm 1, we then present MGDA-FA in Algorithm 4 (shown in Appendix B), where $x_{t}$ is updated along the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to get $x_{t+1}$ following by the update rule of $w_{t}$
@@ 4.1 Deterministic setting
These assumptions are the most relaxed ones in existing MOO works since they directly assume objective smoothness or gradient/function value boundness [25, 13, 29, 35, 6, 36, 3]. It also includes the widely studied standard $L$-smoothness [30, 15, 16], $(L_{0},L_{1})$-smoothness [38] as special cases. Moreover, for any $0\leq\gamma\leq 2$ and $x\in\mathcal{X}$, our assumption even holds for function $f$ such that $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|^{\gamma}$, where $\gamma$ are limited to $[0,1]$ in [9].
We then provide our theoretical results. Let $c>0$ and $F>0$ be some constants such that $\Delta+c\leq F,$ where $\Delta=\max_{i\in[K]}\{f_{i}(x_{0})-f^{*}\}$. Define $M=\sup\{z\geq 0|\varphi(z)\leq F\}$. We then have the following convergence rate for Algorithm 1:
@@
Let Assumptions 1 and 2 hold. Set $\beta\leq\mathcal{O}(\frac{1}{M^{2}}),\alpha\leq\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T\geq\max\Big{(}\Theta\big{(}\frac{1}{\alpha\epsilon^{2}}\big{)},\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\Big{)}$ and $\rho=\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and detailed proof can be found in Appendix C.1. Theorem 1 provides the first convergence rate to obtain an $\epsilon$-accurate Pareto stationary point for MOO problems with generalized $\ell$-smooth objectives. Moreover, it achieves the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-2})$for GD with a single standard $L$-smooth objective [5]. The MOO problems with generalized $\ell$-smooth objectives are challenging due to two reasons: 1) $\|\nabla F(x)\|$ is potentially unbounded in our generalized $\ell$-smoothness setting, making all existing analysis in MOO [25, 13, 29, 35, 6, 36, 3] not applicable. 2) the update of $x$ includes all gradient information from each task, making the existing adaptive methods for single generalized smooth functions invalid.
@@
Under the same setting in Theorem 1, $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}\leq\mathcal{O}(\epsilon^{2})$.
The proof is available in Appendix C.2. Corollary 1 shows that the average CA distance converges.
@@ 4.2 Stochastic setting
Assumption 3 indicates bounded gradient variances, which is widely studied [35, 21, 13].
Let $s_{t,i}=(s_{t,i,1},s_{t,i,2},...,s_{t,i,k})$ be the $i$-th collection of samples at time $t$ in the stochastic model and $F(x_{t};s_{t,i})=(f_{1}(x_{t};s_{t,i,1}),f_{2}(x_{t};s_{t,i,2}),...,f_{k}(x_{t};s_{t,i,k})).$ In this section, we choose $G_{i}(x_{t})=F(x_{t};s_{t,i})$ for any $i\in[3]$. Define $\varepsilon_{t,i}=(\varepsilon_{t,i,1},\varepsilon_{t,i,2},...,\varepsilon_{t,i,k})=\nabla F(x_{t})-\nabla G_{i}(x_{t})$.
@@
Let Assumptions 1, 2, 3 hold. Set $\alpha\leq\min\{\mathcal{O}(\ell(M+1)),\mathcal{O}(\beta),\mathcal{O}(\rho),\mathcal{O}\left(\frac{1}{\sqrt{T}}\right),\mathcal{O}(\epsilon^{2})\}$, $\beta\leq\min\{\mathcal{O}\left(\frac{1}{\alpha{T}}\right),\mathcal{O}(\epsilon^{2}),\mathcal{O}(\rho)\}$, $\rho=\min\{\mathcal{O}\left(\frac{1}{\alpha T}\right),\mathcal{O}(\frac{1}{\sqrt{\alpha T}}),\mathcal{O}\left(\frac{\epsilon}{\sqrt{\beta}}\right),\mathcal{O}(\epsilon^{2})\}$ and $T\geq\Theta\left(\frac{1}{\alpha\epsilon^{2}}+\frac{1}{\epsilon^{4}}\right)$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2},$ with the probability at least $1-\delta$.
The full version and detailed proof can be found in Appendix C.3. When we set $\alpha,\beta,\rho=\mathcal{O}(\epsilon^{2})$ and $T\geq\mathcal{O}(\epsilon^{-4})$, we can find an $\epsilon$-stationary point with the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-4})$ for SGD with a single $L$-smooth objective [1]. Note that in the proof of Theorem 1, we show for each $t\leq T$ and $w\in W$, we have that $F(x_{t})w$ is bounded by applying a small constant step size $\alpha$ and $\beta$. However, this condition does not necessarily hold for our stochastic setting due to the unbounded gradient noise. To solve this problem, we introduce stopping time $\tau$. The advantages are as follows: 1) for any $t\leq\tau$, $w\in W$, we have that $F(x_{t})w$ is bounded; 2) for any $t<\tau$, the norm of gradient noise is bounded; 3) due to the optimal stopping theorem, for any $w\in W$ and $i\in[3]$, we have that $\mathbb{E}[\sum_{t=0}^{\tau}\varepsilon_{t,i}w]=0$. Based on these properties, we can further get the following lemma:
@@
Using the parameters selected in Theorem 2, we have that
$\mathbb{E}[F(x_{\tau})w]-F^{*}w\leq\frac{\delta F}{8}-\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right].$ (7)
The proof of Lemma 1 is available in C.4. Lemma 1 indicates that $\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right]$ is bounded by some constant and if $\tau=T$ with high probability, we have that $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)$. Note that $\{\tau<T\}=\{\tau_{2}<T\}\cup\{\tau_{3}<T\}\cup\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\}.$ The first two events are related to the gradient noise, where the probabilities can be bounded by Assumption 3 and Chebyshev’s inequality. The last event indicates that for some $i\in[K]$, we have $f_{i}(x_{\tau})-f_{i}^{*}\leq\frac{F}{2}.$ Based on Lemma 1 and Markov inequality, we can show that ($\mathbb{P}(\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\})\leq\frac{\delta}{4}$ and we can further show that $\mathbb{P}(\tau=T)\geq 1-\frac{\delta}{2}$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}$ converges with high probability. Similar to Corollary 1, Theorem 2 also implies the average of CA distances converges with time with high probability.

### v5

@@ 2.2 Pareto concepts in multi-objective optimization (MOO)
MOO aims to find points at which there is no common descent direction for all objectives. Considering points $x_{1},x_{2}\in\mathbb{R}^{m}$, we claim that $x_{1}$ dominates $x_{2}$ if $f_{i}(x_{1})\geq f_{i}(x_{2})$ for all $i\in[K]$ and $F(x_{1})\neq F(x_{2})$. We say a point is Pareto optimal if it is not dominated by any other point. In other words, we cannot improve one objective without compromising another when we reach a Pareto optimal point. In non-convex settings, MOO aims to find a Pareto stationary point defined as follows.
@@
$x\in\mathbb{R}^{m}$ is a Pareto stationary point if $\min_{w\in\mathcal{W}}\|\nabla F(x)w\|^{2}=0$, where $\mathcal{W}$ is the probability simplex over $[K]$. $x$ an $\epsilon$-accurate Pareto stationary point if $\min_{w\in\mathcal{W}}\|\nabla F(x)w\|^{2}\leq\epsilon^{2}$.
@@ 2.3 Existing MOO Algorithms
Deterministic MOO. One of the big challenges of MOO is the gradient conflict, i.e., the gradients of different objectives may vary heavily in scale such that the largest gradient dominates the update direction. As a result, the performance of those objectives with smaller gradients (Yu et al., 2020) may be significantly compromised. As the most fundamental MOO algorithm, MGDA tends to find a balanced update direction for all objectives by considering the minimum improvement across all objectives and maximizes it by solving the following problem
@@
where the regularization term $-\frac{1}{2}\|d\|^{2}$ is to regulate the magnitude of our update direction. The solution to the problem in eq. 3 can be obtained by solving the following problem (Désidéri, 2012):
@@
The above approach has been widely used in various variants of MGDA such as SDMGrad, CAGrad, and PCGrad (Xiao et al., 2024; Yu et al., 2020; Liu et al., 2021).
Stochastic MOO. SMG (Liu & Vicente, 2021) is the first stochastic MGDA. It directly replaces the gradients with stochastic gradients and the update rule becomes
@@
where $\nabla F(x;s)$ is the estimate of $\nabla F(x)$ based on the sample $s$. However, this leads to a biased gradient estimation of the update direction $d_{s}^{*}$, and thus it requires an increasing batch size. To solve this issue, another work MoCo (Fernando et al., 2022) introduces a tracking variable $Y$ as a stochastic estimation of the true gradient. Afterward, a double-sampling strategy is proposed by Chen et al. (2024); Xiao et al. (2024) to generate a near-unbiased update direction.
Strong assumptions in analysis. All the works mentioned above require bounded gradients such as Chen et al. (2024); Fernando et al. (2022); Xiao et al. (2024) or $L$-smoothness such as Ban & Ji (2024); Liu et al. (2021); Navon et al. (2022); Yang et al. (2024). Their analyses do not apply to generalized $\ell$-smoothness objectives, since the Lipschitz constant is potentially infinity.
@@ 2.4 Conflict-avoidant (CA) direction and CA distance
The larger the CA distance is, the further the estimated update direction will be away from the CA direction, and the more conflict there will be. In single-loop algorithms, MoCo (Fernando et al., 2022) ensures the average CA distance over iteration is of the order of $\epsilon$, while MoDo (Chen et al., 2024) guarantees the $\epsilon$-order iteration-wise CA weight distance (as stated in their Theorem 3.4). Meanwhile, the double-loop algorithm SDMGrad (Xiao et al., 2024) guarantees an $\epsilon$-order CA distance in every iteration. In this work, we analyze the CA distance in both cases and provide convergence results.
@@ 3.1 MGDA with and without Warm Start
Besides the benefit of a unique solution, adding an $\ell_{2}$ regularization term also makes $w_{\rho}^{*}(x)$ Lipschitz continuous (Fernando et al., 2022). Note that $w^{*}(x)$ may not be Lipschitz continuous because $\nabla F(x)^{\top}\nabla F(x)$ may not be positive definite. Nevertheless, the analysis of CA distance is difficult because $w^{*}$ may not be Lipschitz continuous. Thus, we will characterize the gap between $w^{*}$ and $w_{\rho}^{*}$ plus the change of $w_{\rho}^{*}$ after adding this $\ell_{2}$ regularization term. As a result, the update rules become Lines 5-6 in Algorithm 1. We first update $w_{t}$ by a projected gradient descent process and compute the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to update model parameters.
@@ 3.2 Stochastic MGDA with Double Sampling
In the stochastic setting, our algorithm keeps the same structure, having a warm start process if we aim to control the CA distance in every iteration. In Algorithm 2, we do the same projected gradient descent without using stochastic gradients. This is because we only need to compute $\nabla F(x_{0})^{\top}\nabla F(x_{0})$ once and reuse it in the whole loop, which does not bring a computational burden. Then in the update loop, we update the weight and model parameters accordingly. We use a double-sampling strategy here to make the weight gradient estimator unbiased (Xiao et al., 2024) such that $d_{t}$ is a near-unbiased multi-gradient $\mathbb{E}[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]=\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\rho w_{t},$ where $\nabla G_{2}(x_{t})$ and $\nabla G_{3}(x_{t})$ are independent and unbiased estimates of $\nabla F(x_{t})$. Similarly, we do not involve a warm start process if we require the average CA distance to be small.
3.3 MGDA with Fast Approximation
It can be seen from Algorithm 1 and Algorithm 3 that MGDA requires $\mathcal{O}(K)$ space and time to compute and store all task gradients at each iteration for updating the weight $w_{t}$. This becomes a drawback when the number of tasks or the model size is large. Motivated by Liu et al. (2024), one solution is to use the Taylor Theorem to approximate the gradient for updating the weight $w_{t}$ as
where $R(x_{t})$ is the remainder term and it takes the order $R(x_{t})=o(\|x_{t}-x_{t+1}\|^{2})$, which can be made sufficiently small by adjusting the step size. By incorporating this fast approximation (FA) in Algorithm 1, we then present MGDA-FA in Algorithm 4 (shown in Appendix C), where $x_{t}$ is updated along the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to get $x_{t+1}$ following by the update rule of $w_{t}$
@@ 4.1 Deterministic setting
These assumptions are the most relaxed ones in existing MOO works since they directly assume objective smoothness or gradient/function value boundness (Liu et al., 2021; Fernando et al., 2022; Navon et al., 2022; Xiao et al., 2024; Chen et al., 2024; Yang et al., 2024; Ban & Ji, 2024). It also includes the widely studied standard $L$-smoothness (Nemirovskij & Yudin, 1983; Ghadimi & Lan, 2013; Ghadimi et al., 2016), $(L_{0},L_{1})$-smoothness (Zhang et al., 2019) as special cases. Moreover, for any $0\leq\gamma\leq 2$ and $x\in\mathcal{X}$, our assumption even holds for function $f$ such that $\|\nabla^{2}f(x)\|\leq L_{0}+L_{1}\|\nabla f(x)\|^{\gamma}$, where $\gamma$ are limited to $[0,1]$ in Chen et al. (2023).
Let $c>0$ and $F>0$ be some constants such that $\Delta+c\leq F,$ where $\Delta=\max_{i\in[K]}\{f_{i}(x_{0})-{f_{i}^{*}}\}$. Define $M=\sup\{z\geq 0|\varphi(z)\leq F\}$. We then have the following convergence rate for Algorithm 1:
@@
Let Assumptions 1 and 2 hold. Set $\beta=\mathcal{O}(\frac{1}{M^{2}}),\alpha=\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T=\max\Big{(}\Theta\big{(}\frac{1}{\alpha\epsilon^{2}}\big{)},\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\Big{)}$ and $\rho=\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and detailed proof can be found in Appendix D.1. Theorem 1 provides the first convergence rate to obtain an $\epsilon$-accurate Pareto stationary point for MOO problems with generalized $\ell$-smooth objectives. Moreover, it achieves the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-2})$for GD with a single standard $L$-smooth objective (Carmon et al., 2020). The MOO problems with generalized $\ell$-smooth objectives are challenging due to two reasons: 1) $\|\nabla F(x)\|$ is potentially unbounded in our generalized $\ell$-smoothness setting, making all existing analysis in MOO (Liu et al., 2021; Fernando et al., 2022; Navon et al., 2022; Xiao et al., 2024; Chen et al., 2024; Yang et al., 2024; Ban & Ji, 2024) not applicable. 2) the update of $x$ includes all gradient information from each task, making the existing adaptive methods for single generalized smooth functions invalid.
@@
Under the same setting in Theorem 1, $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}=\mathcal{O}(\epsilon^{2})$.
The proof is available in Appendix D.2. Corollary 1 shows that the average CA distance converges.
@@ 4.2 Stochastic setting
Assumption 3 indicates bounded gradient variances, which is widely studied (Xiao et al., 2024; Li et al., 2024a; Fernando et al., 2022). At time $t$ in the stochastic model, let $s_{t,i}=(s_{t,i,1},s_{t,i,2},...,s_{t,i,k})$ be the $i$-th collection of samples and $F(x_{t};s_{t,i})=(f_{1}(x_{t};s_{t,i,1}),f_{2}(x_{t};s_{t,i,2}),...,f_{k}(x_{t};s_{t,i,k})).$ Note that $i\in[3]$ because we have 3 samples in Algorithm 3. We choose $G_{i}(x_{t})=F(x_{t};s_{t,i})$. Define $\varepsilon_{t,i}=(\varepsilon_{t,i,1},\varepsilon_{t,i,2},...,\varepsilon_{t,i,k})=\nabla F(x_{t})-\nabla G_{i}(x_{t})$.
@@
Let Assumptions 1, 2, 3 hold. Set $\rho=\mathcal{O}(\delta^{2}\epsilon^{2}),\beta=\mathcal{O}(\min\{\frac{\delta\epsilon^{2}}{M^{4}},\frac{\rho}{M^{2}}\}),\alpha=\mathcal{O}(\min\{\beta,\frac{\rho}{\ell(M+1)^{2}},\frac{\rho}{M^{2}},\frac{\delta}{\rho T},\frac{1}{\beta TM^{2}}\})$ and $T=\Theta(\max\{\frac{1}{\delta\alpha\epsilon^{2}},\frac{M^{2}}{\delta^{2}\epsilon^{4}}\})$. We then have that with the probability at least $1-\delta$, $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2})$
The full version and detailed proof can be found in Appendix D.3. When we set $\alpha,\beta,\rho=\mathcal{O}(\epsilon^{2})$ and $T=\Theta(\epsilon^{-4})$, we can find an $\epsilon$-stationary point with the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-4})$ for SGD with a single $L$-smooth objective (Arjevani et al., 2023). Note that in the proof of Theorem 1, we show for each $t\leq T$ and $w\in W$, we have that $F(x_{t})w$ is bounded by applying a small constant step size $\alpha$ and $\beta$. However, this condition does not necessarily hold for our stochastic setting due to the unbounded gradient noise. To solve this problem, we introduce stopping time $\tau$. The advantages are as follows: 1) for any $t\leq\tau$, $w\in W$, we have that $F(x_{t})w$ is bounded; 2) for any $t<\tau$, the norm of gradient noise is bounded; 3) due to the optional stopping theorem, for any $w\in W$ and $i\in[3]$, we have that $\mathbb{E}[\sum_{t=0}^{\tau}\varepsilon_{t,i}w]=0$. Thus, we can further get the following lemma:
@@
Using the parameters selected in Theorem 2, we have that $\mathbb{E}[F(x_{\tau})w]-F^{*}w\leq\frac{\delta F}{8}-\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right].$
The proof of Lemma 1 is available in D.4. Lemma 1 indicates that $\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right]$ is bounded by some constant and if $\tau=T$ with high probability, we have that $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]=\mathcal{O}\left(\frac{1}{\alpha T}\right)$. Note that $\{\tau<T\}=\{\tau_{2}<T\}\cup\{\tau_{3}<T\}\cup\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\}.$ The first two events are related to the gradient noise, where the probabilities can be bounded by Assumption 3 and Chebyshev’s inequality. The last event indicates that for some $i\in[K]$, we have $f_{i}(x_{\tau})-f_{i}^{*}\leq\frac{F}{2}.$ Based on Lemma 1 and Markov inequality, we can show that ($\mathbb{P}(\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\})\leq\frac{\delta}{4}$ and we can further show that $\mathbb{P}(\tau=T)\geq 1-\frac{\delta}{2}$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}$ converges with high probability. Similar to Corollary 1, Theorem 2 also implies the average of CA distances converges with time with high probability.